1. A lawnmower does 6.9 MJ of work in an hour
A. What's its power output, in watts?
B. What's its power output, in horsepower?

Your mass is 65 kg, and you're in an airplane where your gravitational potential energy is7.0 MJ, where the zero of potential is taken at sea level.
What's the plane's altitude?
Answer #1

1) A)

P = Work (Joules) / second, thus,

P = 6.9 * 106 / 3600

P = 1916.67 W.

B) P(hp) = P(W) / 745.699872

P(hp) = 2.57 hp.

2) G.P.E = mgh

7000,000 = 65*9.8*h

h = 10.98 km.